I have worked on this problem for three days, and need some help please.
The question ask: An investment of $57,000 was made by a business club. The investment was split into three parts and lasted for a year.The first part of the investment earned 8% interest, the second 6%, and the third 9% interest. Total interest from the investments was $4290. The interest from the first investment was three times the interest from the second. Find the amounts of the three part investments.
Investment 1 =
Investment 2 =
Investment 3 =
Thanks so much, hopefully you can help!

Answer by stanbon(75887)

An investment of $57,000 was made by a business club. The investment was split into three parts and lasted for a year.The first part of the investment earned 8% interest, the second 6%, and the third 9% interest. Total interest from the investments was $4290. The interest from the first investment was three times the interest from the second. Find the amounts of the three part investments.
-------------
3 equations with 3 variables:
Quantity: x + y + z = 57000
Interest:0.08x+0.06y+0.09z = 4290
Interest:0.08x - 3*0.06y = 0
-----
Rearrange:
x + y + z = 57000
8x + 6y + 9z = 429000
8x - 18y + 0 = 0
=================================
Solve the system by any method you know to get:
x = $36000
y = $16000
z = $5000
=============
Cheers,
Stan H.